# Integration of TinyML and LargeML: A Survey of 6G and Beyond

### v1

@@ IV-D1 Vanilla Federated Split Learning
FSL parallelizes client-side model training (see Section IV-C2) while addressing highly non-IID data across clients, combining the strengths of SL and FL [146]. FSL differs from PSL, EdgeSplit, and EPSL methods by the division of responsibilities between two cooperating servers: the main server (typically a high-performance cloud server) and the Fed server (usually an edge server running FedAvg). In FSL, the Fed server aggregates updated local models from clients using FL-based techniques. This process is illustrated in Fig. 9, including the key steps:
@@
    Simultaneous client-side local FP : Each TinyML device ${\cal N}=\left\{{1,\ldots,N}\right\}$ performs local FP on its model segment, ${\bf{T}}_{n}$, using its local dataset ${{\cal D}_{n}}=\left\{{1,\ldots,{D_{n}}}\right\}$ to generate smashed data ${\bf{T}}_{n}({\bf{x}}_{n})$ at the split layer.
@@
    Main-server processing: The main server processes the smashed data in both local FP and BP using either strategies of SplitFedV1 (when the server has substantial computing capacity and can efficiently execute parallel processing) or SplitFedV2 (personalized updates on the main server are essential, and client data is relatively homogeneous) [146, 147]. In SplitFedV1, each client’s smashed data is separately processed and parallelized on its model segment $\bf{L}$, where the loss value ¯$\bar{\cal L}({{{\bf{y}}_{n}},{{{\bf{\hat{y}}}}_{n}}})$ and its gradient ¯$\nabla\bar{\cal L}({{{\bf{y}}_{n}},{{{\bf{\hat{y}}}}_{n}}})$ are calculated based on the predicted output ${{\bf{\hat{y}}}_{n}}={\bf{L}}({{{\bf{T}}_{n}}({{{\bf{x}}_{n}}})})$ and ground-truth label ${{\bf{y}}_{n}}$. In the subsequent steps, the server aggregates the gradients of the smashed data and updates the server-side model using the FedAvg algorithm [106], expressed as ${\bf{L}}\leftarrow{\bf{L}}-{\eta_{S}}{{\sum\nolimits_{n\in{\cal N}}{{D_{n}}\nabla\bar{\cal L}\left({{{\bf{y}}_{n}},{{{\bf{\hat{y}}}}_{n}}}\right)}}}\Big{/}{{\sum\nolimits_{n\in{\cal N}}{{D_{n}}}}}$. In SplitFedV2, the server processes client smashed data sequentially on $\bf{L}$, following a randomly determined client order represented by the list $\cal S$ and the computation of ¯$\nabla\bar{\cal L}\left({{{\bf{y}}_{n}},{{{\bf{\hat{y}}}}_{n}}}\right)$ akin to SplitFedV1. Instead of aggregation, the server sequentially process the data in the order defined by $\cal S$ to update the global model after each local FP using the rule ${\bf{L}}\leftarrow{\bf{L}}-{\eta_{S}}\nabla\bar{\cal L}\left({{{\bf{y}}_{n}},{{{\bf{\hat{y}}}}_{n}}}\right)$ and continuing until all clients in the list $\cal S$ are processed.
@@
    BP of smashed data gradients: After performing local FP and BP on the server-side model, the main server returns the evaluation of ¯$\nabla\bar{\cal L}\left({{{\bf{y}}_{n}},{{{\bf{\hat{y}}}}_{n}}}\right)$ to TinyML devices.
@@
    Client-side local BP: Upon receiving the gradients from the main server, each TinyML device $n\in{\cal N}$ calculates ¯$\nabla\bar{\cal L}\left({{{\bf{T}}_{n}}\left({{{\bf{x}}_{n}}}\right)}\right)$ and performs local BP, updating its model as ${{\bf{T}}_{n}}\leftarrow{{\bf{T}}_{n}}-{\eta_{n}}\nabla\bar{\cal L}\left({{{\bf{T}}_{n}}\left({{{\bf{x}}_{n}}}\right)}\right)$.
@@
Several recent studies have explored the use of FSL for integrating TinyML and LargeML in a range of applications [148, 149, 150, 151]. For example, AdaptSFL, based on SplitFedV2, dynamically adjusts model splitting and aggregation to optimize latency and convergence on resource-constrained devices, outperforming benchmarks [148]. FedSL, also on SplitFedV2, enables robust healthcare analytics on wearables, excelling in medical imaging across diverse data [149]. Beyond standard FSL architectures,SplitGP, built on SplitFedV1, supports bidirectional training for personalized client models and generalized server models, surpassing baselines [150]. ESFL, also on SplitFedV1, enhances training stability by incorporating prior global models, using an iterative algorithm to solve its NP-hard optimization, improving efficiency over standard FL, SL, and FSL [151]. Hierarchical FSL frameworks further boost efficiency: one trains multiple edge server-client pairs with cloud aggregation for resilience [152], while another uses parallel multi-client setups for enhanced performance on constrained IoT devices [153].
IV-D2 Federated Split Foundation Models and Federated Split Fine-Tuning of Foundation Models
@@
Federated split FMs (FSFMs) extend the FL framework by partitioning large FMs across clients and servers, enabling distributed training and inference with reduced local resource requirements [154]. Clients train or fine-tune distinct FM segments collaboratively, preserving privacy by not exposing raw data, and lowering computational overhead. Federated split fine-tuning (FedSFT) further enables client-side personalization of model components while retaining global knowledge through coordinated updates. Compared to standard FL and FFMs, FSFMs improve scalability by exchanging only compact smashed data rather than full model parameters [154]. These features position FSFMs and FedSFT as particularly suitable approaches in 6G scenarios characterized by device heterogeneity, limited computational capacity, and stringent privacy requirements.
@@
For instance, SFPrompt is a prompt-based FedSFT framework designed for privacy-sensitive, resource-constrained edge environments [155]. By dividing the pre-trained FM between server and client and integrating soft prompts with local data pruning, SFPrompt improves FedFT efficiency and performance and significantly reduces computational and communication costs in image classification tasks. FedSplitX is a fine-grained FSFM approach that supports multiple model partition points to accommodate diverse client capabilities in heterogeneous, resource-constrained edge environments [156]. By embedding auxiliary networks at each partition, FedSplitX reduces communication overhead and latency and improves model performance for image classification tasks. FedVZ provides a defense framework against gradient inversion attacks in FSFMs using vision transformers [157]. It replaces BP with a single forward pass based on zeroth-order optimization to approximate server gradients, proving effective in heterogeneous, resource-limited environments for image classification tasks. PRINCE provides an incentive mechanism that encourages high-quality device participation from multiple FSL tenants through strategic pricing [158]. PRINCE improves FedSFT performance across a range oftasks, including image classification, sentiment analysis, speed-to-text, and question answering.
@@ V-A Data Privacy and Network Security
V-A1 Toward Data Privacy
Model inversion attacks present significant privacy threats, particularly for systems trained on sensitive data. TinyML–LargeML integration enables a robust privacy-preserving architecture through on-device inference and decentralized data processing. TinyML enhances data privacy by performing local processing on edge devices [11], minimizing dependence on centralized cloud servers and reducing the exposure of raw data. In parallel, LargeML complements this architecture by aggregating and analyzing processed data without accessing the original inputs [159]. By sharing only abstracted features or model updates, a collaborative model mitigates the risks associated with model inversion attacks. TinyML also facilitates the deployment of conventional privacy-preserving techniques such as FL, allowing distributed devices to train models collaboratively without sharing raw data [11]. Differential privacy can further enhance protection by injecting noise into data or model outputs, preventing identification of individual users while maintaining meaningful analysis [160]. When combined with LargeML, more advanced FL-based techniques, such as FTL [104], FML [115], and FSL [146], enable stronger privacy protection in AI ecosystems, as detailed in Section IV.
Privacy-preserving integration is particularly crucial in healthcare applications [99, 109, 119, 149], where wearable devices embedded with TinyML algorithms can process physiological data locally. Personal health information thus remains private, while LargeML systems, hosted on secure servers, draw from aggregated insights to improve diagnostic accuracy and predictive performance without compromising raw data. Beyond healthcare, TinyML–LargeML integration benefits a range of privacy-sensitive domains. For example, in smart agriculture [101], IoT-enabled agricultural sensors equipped with TinyML analyze soil conditions, crop health, and environmental metrics on-site, protecting proprietary farming practices, while LargeML refines predictive models for climate adaptation and yield forecasting. In smart home and city [118], on-device inference captures user behavior and preferences with minimal data exposure, while LargeML enhances service personalization through cloud-based analysis.
V-A2 Toward Network Security
As reviewed in Section III-A, TinyML devices operate under stringent resource constraints, with memory and computing capacity typically two to three orders of magnitude below that of typical IoT or edge devices [161, 56, 11]. These limitations make traditional security mechanisms inadequate, and integrating TinyML and LargeML is projected to introduce a novel secure strategy. In particular, TinyML enables lightweight, real-time anomaly detection on edge devices, identifying threats with low computational cost by monitoring local activity for deviations [162]. Integrated with LargeML, which analyzes large-scale network traffic and correlates anomalies across devices [163], this forms a multi-layered defense for 6G networks. TinyML handles local threat detection, while LargeML uncovers complex attack patterns. For instance, IoTDefender [113] uses FTL across three edge servers to aggregate local models, refined by a cloud-based CNN for pre-training. During TL, models are adapted for IoT devices, enabling efficient intrusion detection despite hardware constraints.
To support efficient encryption and secure communication, TinyML enhances data security by performing lightweight encryption and decryption on edge devices, enabling secure data transmission without relying solely on cloud services [164]. However, its limited resources restrict complex encryption protocols. LargeML addresses this by managing encryption key distribution and security protocol orchestration, dynamically updating keys to strengthen communication channels [165]. This hybrid approach balances robust security with TinyML’s constraints. Fig. 10 depicts a TinyML–LargeML ecosystem enhancing 6G data privacy and network security.
@@ V-B Network Management
With the rapid growth of connected IoT devices and mobile users, 5G networks have been engineered to manage high traffic volumes and support ultra-reliable low-latency communication (uRLLC), massive machine type communication (mMTC), and enhanced mobile broadband (eMBB) using technologies such as network slicing, software-defined networking (SDN), network function virtualization (NFV), and multi-access edge computing (MEC) [166]. In which, network slicing creates virtual networks on shared infrastructure, SDN and NFV enhance flexibility by decoupling control/data planes and virtualizing functions, and MEC reduces latency by processing data at the edge. As 6G emerges, user-centric, intelligent management increases complexity, rendering traditional approaches inadequate. Closed-loop automation, powered by ML and big data analytics, is now critical [167].
Against this background, TinyML–LargeML integration transforms 6G network management by combining edge-based efficiency with large-scale analytics for adaptive, intelligent systems. TinyML enables localized optimization, monitoring real-time metrics like latency and traffic to ensure low-latency and quality-of-service (QoS) in applications like smart factories [168]. LargeML processes data from multiple TinyML nodes, identifying complex patterns, predicting failures, and optimizing network-wide resources, such as rerouting traffic to maintain uRLLC, mMTC, and eMBB connectivity [169]. For instance, a Lyapunov-based adaptive split learning framework dynamically assigns split layers to TinyML devices, reducing latency and energy use [132]. Fig. 11 is an illustrative combination of TinyML’s efficiency at the edge with LargeML’s capacity for large-scale analytics, and applications of bidirectional integration across domains. In network slicing, TinyML monitors metrics for real-time resource reallocation, while LargeML predicts congestion and optimizes slice provisioning [170, 171]. In SDN, TinyML detects anomalies like DDoS attacks at the edge, and LargeML enhances traffic management by predicting congestion and optimizing routes [172, 173]. For NFV, TinyML addresses performance degradation, while LargeML automates resource scaling [174]. In MEC, TinyML reduces latency through on-device processing, complemented by LargeML’s orchestration, ensuring efficient, low-latency 6G networks [175]. For example, SplitGP [150], an FSL strategy combining personalization and generalization for training-inference optimizations. Another is SplitGP, which optimizes the TinyML device-side model for task-specific personalization and trains the MEC-side model to generalize across out-of-distribution inputs.

### v2

@@ IV-D1 Vanilla Federated Split Learning
Vanilla FSL, the simplest and foundational form of FSL, merges the core ideas of SL, where the model is cut at a chosen layer and clients only compute the early part of the network, with FL, where many clients participate in training while keeping their data local. This process is illustrated in Fig. 9, with the key steps:
@@
    Simultaneous client-side local FP: Each TinyML device $n\in{\cal N}$ performs local FP on its model segment, ${\bf{T}}_{n}$, using its local dataset ${\cal D}_{n}$} to generate smashed data ${\bf{T}}_{n}({\bf{x}}_{n})$ at the split layer.
@@
    Main-server processing: The main server processes the smashed data in both local FP and BP using either strategies of SplitFedV1 (when the server has substantial computing capacity and can efficiently execute parallel processing) or SplitFedV2 (personalized updates on the main server are essential, and client data is relatively homogeneous) [159, 158]. In SplitFedV1, each client’s smashed data is separately processed and parallelized on its model segment $\bf{L}$, where the loss value $\bar{\cal L}({{{\bf{y}}_{n}},{{{\bf{\hat{y}}}}_{n}}})$ and its gradient $\nabla\bar{\cal L}({{{\bf{y}}_{n}},{{{\bf{\hat{y}}}}_{n}}})$ are calculated based on the predicted output ${{\bf{\hat{y}}}_{n}}={\bf{L}}({{{\bf{T}}_{n}}({{{\bf{x}}_{n}}})})$ and ground-truth label ${\bf{y}}_{n}$}. In the subsequent steps, the server aggregates the gradients of the smashed data and updates the server-side model using the FedAvg algorithm [113], expressed as ${\bf{L}}\leftarrow{\bf{L}}-{\eta_{S}}{{\sum\nolimits_{n\in{\cal N}}{{|{\cal D}_{n}|}\nabla\bar{\cal L}\left({{{\bf{y}}_{n}},{{{\bf{\hat{y}}}}_{n}}}\right)}}}\Big/{{\sum\nolimits_{n\in{\cal N}}{{|{\cal D}_{n}|}}}}$. In SplitFedV2, the server processes client smashed data sequentially on $\bf{L}$, following a randomly determined client order represented by the list $\cal S$ and the computation of $\nabla\bar{\cal L}\left({{{\bf{y}}_{n}},{{{\bf{\hat{y}}}}_{n}}}\right)$ akin to SplitFedV1. Instead of aggregation, the server sequentially process the data in the order defined by $\cal S$ to update the global model after each local FP using the rule ${\bf{L}}\leftarrow{\bf{L}}-{\eta_{S}}\nabla\bar{\cal L}\left({{{\bf{y}}_{n}},{{{\bf{\hat{y}}}}_{n}}}\right)$ and continuing until all clients in the list $\cal S$ are processed.
@@
    BP of smashed data gradients: After performing local FP and BP on the server-side model, the main server returns the evaluation of $\nabla\bar{\cal L}\left({{{\bf{y}}_{n}},{{{\bf{\hat{y}}}}_{n}}}\right)$ to TinyML devices.
@@
    Client-side local BP: Upon receiving the gradients from the main server, each TinyML device $n\in{\cal N}$ calculates $\nabla\bar{\cal L}\left({{{\bf{T}}_{n}}\left({{{\bf{x}}_{n}}}\right)}\right)$ and performs local BP, updating its model as ${{\bf{T}}_{n}}\leftarrow{{\bf{T}}_{n}}-{\eta_{n}}\nabla\bar{\cal L}\left({{{\bf{T}}_{n}}\left({{{\bf{x}}_{n}}}\right)}\right)$.
@@
Several recent studies have explored the use of FSL for integrating TinyML and LargeML in a range of applications [55, 210, 100, 123]. Building upon SplitFedV1, SplitGP supports bidirectional training for personalized client models and generalized server models, surpassing baselines [55]. Similarly, ESFL enhances training stability by incorporating prior global models, using an iterative algorithm to solve its NP-hard optimization, improving efficiency over standard FL, SL, and FSL [210]. For SplitFedV2, AdaptSFL dynamically adjusts model splitting and aggregation to optimize latency and convergence on resource-constrained devices, outperforming benchmarks [100]. Meanwhile, FedSL enables robust healthcare analytics on wearables, excelling in medical imaging across diverse data [123]. Further, hierarchical FSL frameworks boost efficiency: one trains multiple edge server-client pairs with cloud aggregation for resilience [206], while another uses parallel multi-client setups for enhanced performance on IoT devices [78].
IV-D2 Federated Split Models and Fine-Tuning
@@
Federated split FMs (FSFMs) extend the FL framework by partitioning large FMs across clients and servers, enabling distributed training and inference with reduced local resource requirements [93]. Clients train or fine-tune distinct FM segments collaboratively, preserving privacy by not exposing raw data and lowering computational overhead. Federated split fine-tuning (FedSFT) further enables client-side personalization of model components while retaining global knowledge through coordinated updates. Compared to standard FL and FFMs, FSFMs improve scalability by exchanging only compact smashed data rather than full model parameters [93]. These features position FSFMs and FedSFT as particularly suitable approaches in 6G scenarios characterized by device heterogeneity, limited computational capacity, and stringent privacy requirements.
@@
For instance, SFPrompt [21] is a prompt-based FedSFT framework for privacy-sensitive, resource-constrained edge environments. By dividing pre-trained FMs between server and client and integrating soft prompts with local data pruning, it improves FedFT efficiency while significantly reducing computation and communication costs in image classification. FedSplitX [151] introduces fine-grained FSFM with multiple model partition points to accommodate diverse client capabilities in heterogeneous, resource-constrained edge environments. Auxiliary networks at each partition reduce latency and communication overhead, enhancing image classification accuracy. FedVZ [150] defends against gradient inversion in FSFMs using vision transformers by replacing BP with a zeroth-order optimization forward pass, enabling secure, efficient training in constrained environments. Finally, PRINCE [94] incentivizes high-quality device participation from multiple FSL tenants via strategic pricing, boosting FedSFT performance in tasks such as image classification, sentiment analysis, speech-to-text, and question answering.
Table IV: Summary of TinyML–LargeML integration solutions.
Sol. Procedures Characteristics Representatives      Advances
TL $(1)$ Server-side training $\to$ $(2)$ Pre-trained knowledge transfer $\to$ $(3)$ Model personalization and on-device adaptation - Aligns well with resource-constrained TinyML devices - Requires ideal conditions: source and target domains with distinct yet closely related datasets and tasks TinyTL [19], [125], [131], [12], [57], [58], [84], and [200] Multi-level server collaboration: Hybrid training, distributed fashion, hierarchical processing, and enhanced availability
FTL $(1)$ Model initialization $\to$ $(2)$ Source-client-side local training $\to$ $(3)$ Local model upload $\to$ $(4)$ Aggregation $\to$ $(5)$ Global model update $\to$ $(6)$ Knowledge transfer $\to$ $(7)$ Model personalization and on-device adaptation (Steps (2)(5)(2)-(5) iterate until convergence) - Enables collaborative learning and distinguishes between source and target clients - Combines the strengths of FL and TL, efficient in highly non-IID data across TinyML devices TinyFedTL [80], FedHealth [26], ACGAN-FTL [54], and [48] - Hierarchical FTL: IoTDefender [43], and HFTL [132] - FML: [141, 140, 142, 73] - FFM and FedFT of FMs: [50, 29, 130, 11, 192, 190, 114, 166]
VSL $(1.1)$ First client local FP $\to$ $(1.2)$ FP $\to$ $(1.3)$ Server-side local FP and BP $\to$ $(1.4)$ BP $\to$ $(1.5)$ First client local BP $\to\cdots\to$ $(N.1)$ Last client local FP $\to\cdots\to$ $(N.5)$ Last client local BP $\to$ $(6)$ On-device adaptation (Steps (1.1)(N.5)(1.1)-(N.5) iterate until convergence) - Enables collaborative learning with sequential training among participating TinyML devices - Incurs high training latency, lacks scalability, overfits to the last trained client ARES [148], frameworks in [88] and [95] Multi-level server collabo-ration: Multi-hop SL [181] and centralized ad-hoc data routing [102]
PSL $(1)$ Simultaneous client-side local FP $\to$ $(2)$ FP $\to$ $(3)$ Server-side local FP and BP $\to$ $(4)$ BP $\to$ $(5)$ Simultaneous client-side local BP $\to$ $(6)$ On-device adaptation (Steps (1)(5)(1)-(5) iterate until convergence) - Enables collaborative learning with parallel training among TinyML devices - Significantly reduces training latency, offers “semi-scalability”, and partially addresses non-IID data issues AdaSplit [30], frameworks in [72] and [79] - Multi-level server collaboration: Parallel-server PSL [162] and multi-hop PSL [161] - Advanced PSL: EdgeSplit [204] and EPSL [101]
Edge Split $(1)$ Simultaneous client-side local FP $\to$ $(2)$ FP1 $\to$ $(3)$ Server-side local FP and BP $\to$ $(4)$ BP1 $\to$ $(5)$ Client-side local BP $\to$ $(6)$ FP2 $\to$ $(7)$ Server-side aggregation $\to$ $(8)$ BP2 $\to$ $(9)$ Updated client-side model $\to$ $(10)$ On-device adaptation (Steps (1)(9)(1)-(9) iterate until convergence) - An advanced variant of PSL with additional server-side aggregation - Improves convergence rate in highly non-IID data environments while maintaining PSL benefits [204] CPSL [191] for clustering multiple groups of TinyML devices
EPSL $(1)$ Simultaneous client-side local FP $\to$ $(2)$ FP $\to$ $(3)$ Server-side local FP $\to$ $(4)$ Last-layer gradient aggregation and BP $\to$ $(5)$ BP1 $\to$ $(6)$ BP2 $\to$ $(7)$ Client-side local BP $\to$ $(8)$ On-device adaptation (Steps (1)(7)(1)-(7) iterate until convergence) - An advanced variant of PSL with modified server-side aggregation - Maintains PSL benefits, reduces computational workload and training latency, offers full scalability [101]
FSL The following procedures are for both SplitFedV1 and SplitFedV2, except that step $(3+)$ is specified to SplitFedV1: $(1)$ Simultaneous client-side local FP $\to$ $(2)$ FP $\to$ $(3)$ Main-server-side local FP $\to$ $(3+)$ Main-server-side aggregation and local BP (only for SplitFedV1) $\to$ $(4)$ BP $\to$ $(5)$ Client-side local BP $\to$ $(6)$ Client-side local model upload $\to$ $(7)$ Fed-server aggregation $\to$ $(8)$ Client-side global model update $\to$ $(9)$ Client-side local update $\to$ $(10)$ On-device adaptation (Steps (1)(9)(1)-(9) iterate until convergence) - SplitFedV1: An advanced variant of PSL, with aggregation on both server and client sides through main and Fed servers; Combines the strengths of FL and SL, improves convergence rate in highly non-IID data environments - SplitFedV2: An advanced variant of PSL with client-side aggregation through Fed server support; Combines the strengths of FL and SL, improves convergence in highly non-IID data environments AdaptSFL [100], FedSL [123], SplitGP [55], and ESFL [210] - Hierarchical FSL: Frameworks in [206] and [78] - FSFMs and FedSFT of FMs: SFPrompt [21], FedSplitX [151], FedVZ [150], and PRINCE [94]
IV-E Discussions and Lessons Learned
While Table IV outlines their key features and advances, lessons learned from the TinyML–LargeML integration provide valuable insights for future research.
First, TL-based integration (Section IV-A) fine-tunes pre-trained models on TinyML devices by updating only a subset of parameters, reducing computational overhead and enabling on-device adaptation. However, it assumes closely related source and target domains, which limits generalizability.
Second, FTL-based integration (Section IV-B) enables client-level personalization under highly non-IID data, aligning well with 6G ubiquitous connectivity. FML further enhances adaptability through meta-learning, while FFMs combine FMs and FL to support collaborative, privacy-preserving, and decentralized FedFT; efficiency is improved via PEFT, prompt tuning, and MoE to handle model heterogeneity.
Third, SL-based integration (Section IV-C) allows resource-constrained TinyML devices to participate in full model training, but VSL suffers from high latency, limited scalability, and overfitting risks due to sequential training. PSL mitigates these issues through parallelization and partial non-IID handling. Building on PSL, EdgeSplit improves convergence under highly non-IID settings, while EPSL further reduces server-side workload and training latency, achieving full scalability
Finally, FSL-based integration (Section IV-D) combines SL and FL to enable flexible aggregation and improved convergence under highly non-IID data, with SplitFedV1/V2 serving as foundational designs. Extensions such as FSFMs and FedFT of split FMs support scalable, privacy-preserving personalization via lightweight smashed-data exchange, though their SL-based architectures remain sensitive to communication latency and network conditions.
Notably, interdisciplinary collaboration is continued to advance and refine TinyML–LargeML integration strategies. Active industry participation is equally important to ensure that proposed solutions are both innovative and practical for real-world deployment. Although 6G has not yet been officially standardized and it remains uncertain whether these approaches will fully satisfy the requirements of 6G and future networks, they represent meaningful progress toward next-generation intelligent wireless networks.
@@ V-A Data Privacy and Network Security
Toward Network Privacy: Model inversion attacks pose major risks when handling sensitive data, and integrating TinyML with LargeML provides a privacy‑preserving architecture by keeping inference and feature extraction on-device while sending only abstracted representations or model updates to the cloud [1, 199]. TinyML supports established techniques such as federated learning [1] and differential privacy [176], while its combination with LargeML enables stronger ecosystem-wide protection through FTL [195], FML [106], and FSL [159] (see Section IV). This integration is especially important in healthcare [131, 26, 73, 123], where TinyML wearables process physiological signals locally and LargeML servers use aggregated insights for diagnosis and prediction; similar benefits appear in smart agriculture [57], where on-site analysis protects proprietary practices while cloud models improve forecasting, and in smart homes and cities [142], where local inference preserves user privacy while LargeML enhances personalization and resource optimization.
Toward Network Security: TinyML devices operate under strict memory and compute limits, often two to three orders of magnitude below typical IoT or edge hardware [98, 96, 1], making conventional security mechanisms impractical, yet their integration with LargeML enables a multi-layered 6G defense in which TinyML performs lightweight, real‑time anomaly detection at the edge [10] while LargeML correlates traffic patterns across devices to uncover complex attacks [200], as demonstrated by IoTDefender [43], which applies FTL across edge servers and cloud CNNs for intrusion detection under hardware constraints. To secure communication, TinyML supports lightweight encryption on-device [129], while LargeML manages key distribution and protocol orchestration, dynamically updating keys to reinforce communication channels [18]. This hybrid approach balances strong security with device limitations and underpins the 6G privacy‑and‑security ecosystem illustrated in Fig. 10.
@@ V-B Network Management
The rise of IoT devices and mobile users has pushed 5G networks to support ultra-reliable low-latency communication (uRLLC), massive machine-type communication (mMTC), and enhanced mobile broadband (eMBB) using network slicing, software-defined networking (SDN), network function virtualization (NFV), and multi-access edge computing (MEC) [16]. Here, network slicing creates virtual networks on shared infrastructure, SDN and NFV provide flexibility through decoupling and virtualization, and MEC reduces latency via edge processing. As 6G moves toward user-centric and intelligent management, these traditional mechanisms become insufficient, necessitating closed-loop automation powered by ML and big data analytics [28].
TinyML–LargeML integration offers a transformative approach by combining edge efficiency with large-scale analytics for adaptive management. TinyML enables localized optimization by monitoring metrics, like latency and traffic, to sustain low-latency quality-of-service (QoS) in applications such as smart factories [153]. LargeML aggregates data across nodes, detects complex patterns, predicts failures, and orchestrates network-wide resources, for instance, rerouting traffic to preserve uRLLC, mMTC, and eMBB [9]. A notable example is a Lyapunov-based adaptive split learning framework that dynamically assigns split layers to TinyML devices, reducing latency and energy consumption [95]. Fig. 11 illustrates such bidirectional integration across domains. Applications span major technologies: (i) In network slicing, TinyML monitors metrics for real-time resource reallocation, while LargeML predicts congestion and optimizes slice provisioning [77, 15]; (ii) In SDN, TinyML detects anomalies (e.g., DDoS attacks) at the edge, and LargeML enhances traffic control by predicting congestion and optimizing routes [85, 7]; (iii) In NFV, TinyML mitigates performance degradation, while LargeML automates resource scaling [120]; (iv) In MEC, TinyML minimizes latency through on-device inference, complemented by LargeML’s orchestration to ensure efficient, low-latency operation [20]. Examples include SplitGP [55], an FSL strategy that balances personalization (on-device TinyML) with generalization (MEC-side LargeML), enabling robust training-inference optimization across diverse inputs.